# Signal Recovery From Product of Two Vandermonde Matrices

### 27 Jan 2024 v1

@@ II-B Derivation of R2
    Compute $\bm{g}=\left(\bm{V}(\bm{z})^{T}\bm{V}(\bm{\theta}_{\bm{p}})\right)^{\dagger}\bm{y}$ or use the procedures outlined in the proofs of Proposition 1 and Proposition 2.
@@ III Phase-less Measurements
  R3: There exists a construction of an $m\times n$ measurement matrix $\bm{A}$, with $m\geq 4s$ and $n\geq 4s-1$, such that there is a unique $\bm{x}\in\mathbb{C}_{\leq s}^{n}$ satisfying
$\bm{y}=\left|\bm{A}\bm{x}\right|^{2}.$ (52)
Furthermore, there exists a corresponding recovery algorithm such that
@@
  This result provides a direct solution to open problem 3.5 in [18], as explained in Section I. To obtain R3, we begin by considering the phase-less equivalent of R1 with a Fourier matrix
@@
  • If $4s-1\leq m<\min\{n+1,8s-3\}$ and samples in $\bm{z}$ are shifted harmonics, $\bm{\theta}$ and $|\bm{g}|$ (up to a global real scalar) can be exactly recovered with $O(\mathrm{poly}(s))$ complexity and there are $2^{s-1}$ different $\bm{g}$ satisfying (53).
  • If $m\geq 8s-3$ and $\bm{z}$ is arbitrary, $\bm{\theta}$ and $|\bm{g}|$ (up to a global real scalar) can be exactly recovered with $O(\mathrm{poly}(s))$ complexity and there are only two different $\bm{g}$ satisfying (53).
To derive R4 we will start (as with R1) by first recovering $\bm{\theta}$ from (53). However, in contrast to R1, we can no longer recover $\bm{g}$ with a simple matrix inversion due to the non-linear magnitude operator. After support recovery $\bm{\theta}$, we are (at first glance) essentially back to a classical phase retrieval problem. Below we will present an approach, similar in nature to the approach in Section II, from which certain Laurent polynomials are solved for. It turns out that there is enough information about $\bm{g}$ embedded in these Laurent polynomials.
R4 shows that there is an ambiguity when measuring as in (53) which cannot be resolved by simply taking more measurements (increasing $m$). In order to resolve the ambiguity of which $\bm{g}$ is the correct one, we take one additional measurement with some arbitrary measurement vector $\bm{a}$. Therefore, we have the following reformulation of R4
@@
  • If $m\geq 8s-3$ and $\bm{z}$ is arbitrary, $\bm{\theta}$ and $\bm{g}$ can be recovered exactly with $O(\mathrm{poly}(s))$ complexity.
@@ III-A Derivation of R4 and R5
where the operator ${.}\mathrm{fliplr}\{.\}$ flips a matrix vertically (i.e., flips its columns), $\bm{G}(\bm{z})$ has dimensions $m\times(8s-2)$ and $\bm{w}=[\hat{\bm{l}};\tilde{\bm{l}};\bm{l};\overline{\tilde{\bm{l}}}]$. Assume that $S$ values from $g_{0},\ldots,g_{s-1}$ are non-zero. Since $\hat{L}(z)=\hat{L}^{+}(z),L(z)=L^{+}(z),\tilde{L}(z)=\tilde{L}^{+}(z)$ solves (60),
@@
solves (III-A) and vice versa. This implies that $\bm{G}(\bm{z})$ cannot be full rank and therefore has a null space of at least one dimension. If $\bm{G}(\bm{z})$ has a one-dimensional null space, meaning that (III-A) has the solution $\bm{w}=c\bm{w}^{+}$ for any real scalar $c$, then (60) has the unique solution $\hat{L}(z)=c\hat{L}^{+}(z),L(z)=cL^{+}(z),\tilde{L}(z)=c\tilde{L}^{+}(z)$ and vice versa.
@@
for any real scalar $c$, which is equivalent to the solution $\hat{L}(z)=c\hat{L}^{+}(z)$, $L(z)+e^{i\gamma/n}\tilde{L}(z)+e^{-i\gamma/n}\overline{\tilde{L}(z)}=c\left(L^{+}(z)+e^{i\gamma/n}\tilde{L}^{+}(z)+e^{-i\gamma/n}\overline{\tilde{L}(z)^{+}}\right)$ of (72) and vice versa.
@@
Assume that $n\geq 4s-1$, $z_{j}^{n}=e^{i\gamma}$ for $20\leq j\leq 4s-$ 2, $g_{k}^{+}\neq 0$, $\theta_{k}^{+}\in\mathbb{T}$ and $(\theta_{k}^{+})^{n}\neq e^{-i\gamma}$ for $10\leq k\leq s-$ 1. Solving (72) (or equivalently (III-A)) results in $\hat{L}(z)=c\hat{L}^{+}(z),L(z)+z^{n}\tilde{L}(z)+z^{-n}\overline{\tilde{L}}(z)=c\left(L^{+}(z)+z^{n}\tilde{L}^{+}(z)+z^{-n}\overline{\tilde{L}^{+}}(z)\right)$ for some real scalar $c$.
@@
Since $(\theta_{k}^{+})^{n}\neq e^{-i\gamma}$ and $g_{k}\neq 0$, it follows that $t_{k}^{+}\left(\frac{1}{\theta_{k}^{+}}\right)\neq 0$ and thus $g_{k}^{+}t_{k}^{+}\left(\frac{1}{\theta_{k}^{+}}\right)\neq 0$, which implies that the roots of $\hat{L}^{+}(z)$ are not roots of $L^{+}(z)+e^{i\gamma}\tilde{L}^{+}(z)+e^{-i\gamma}\overline{\tilde{L}^{+}(z)}$, a contradiction. Hence, the roots of $\hat{L}^{+}(z)$ are roots of $\hat{L}(z)$ which implies that $\hat{L}(z)=c\hat{L}^{+}(z)$ for some real scalar $c$.
@@
From the result in Theorem 7, it follows that when $g_{k}\neq 0$, $10\leq k\leq s-$ 1, we can find $\hat{L}^{+}(z)$ (up to a real scalar $c$) after $14s-$ 1 measurements with shifted harmonic samples from which we can recover the support $\bm{\theta}^{+}$ (below we will deal with the case when $S$, $S\leq s$, of $g_{0},\ldots,g_{s-1}$ are non-zero). Furthermore, from these measurements, we also obtain $L^{+}(z)+e^{i\gamma}\tilde{L}^{+}(z)+e^{-i\gamma}\overline{\tilde{L}^{+}(z)}=\left|e^{i\gamma}\hat{u}^{+}(z)+\tilde{u}^{+}(z)\right|^{2}$ (up to real scalar $c$). For arbitrary $\bm{\theta}^{+}\in\mathbb{T}^{s}$ and $\bm{g}^{+}$, the assumptions of Theorem 7 are satisfied w.p.1, implying that $\tilde{\bm{G}}(\bm{z})$ in (III-A) has a one-dimensional null space w.p.1 when the $m\geq 4s-1$ samples in $\bm{z}$ are shifted harmonics. Moreover, for $\bm{g}^{+}$ with no zero elements and any $\bm{\theta}^{+}\in\mathbb{T}^{s}$, it is evidently easy to choose a $\bm{z}$ such that the assumptions of Theorem 7 are satisfied.
From Theorem 7 one can find $|\bm{g}^{+}|$ up to a real scalar.
Proposition 5.
Given the assumptions in Theorem 7, $|\bm{g}^{+}|$ can be found, up to a real scalar $c$, with $O(\mathrm{poly}(s))$ complexity.
@@
for some unknown real scalar $c$. Since $(\theta_{k}^{+})^{-1}=\overline{\theta_{k}^{+}}$ and
@@
Clearly, the operations leading to $c|\bm{g}^{+}|$ are of $O(\mathrm{poly}(s))$ complexity, showing that $c|\bm{g}^{+}|$ can be found with $O(\mathrm{poly}(s))$ complexity. ∎
@@
When it comes to recovering $\bm{g}^{+}$, the situation is trickier. Given $\left|e^{i\gamma}\hat{u}^{+}(z)+\tilde{u}^{+}(z)\right|^{2}$, up to a real scalar, Proposition 4 implies that the roots of $\left|e^{i\gamma}\hat{u}^{+}(z)+\tilde{u}^{+}(z)\right|^{2}$ appear as conjugate pairs: if $r$ is a root of $\left|e^{i\gamma}\hat{u}^{+}(z)+\tilde{u}^{+}(z)\right|^{2}$ then so is $1/\overline{r}$ as well. Let $(r_{0},1/\overline{r_{0}}),\ldots,(r_{s-2},1/\overline{r_{s-2}})$ be the $s-1$ conjugate root pairs. As seen from (II-A), $e^{i\gamma}\hat{u}^{+}(z)+\tilde{u}^{+}(z)$ is of the same form as $\tilde{u}^{+}(z)$ in (II-A) but with coefficients $(e^{i\gamma}(\theta_{k}^{+})^{n}-1)g_{k}^{+}$, $10\leq k\leq s-$ 1. Hence, for an arbitrary $\bm{g}^{+}$, Lemma 2 implies that $(r_{0},1/\overline{r_{0}}),\ldots,(r_{s-2},1/\overline{r_{s-2}})$ are distinct w.p.1. Only one member from each pair can be a root of $e^{i\gamma}\hat{u}^{+}(z)+\tilde{u}^{+}(z)$. Since we do not know beforehand which member of each pair is the actual root of $e^{i\gamma}\hat{u}^{+}(z)+\tilde{u}^{+}(z)$, we need to go through all the $2^{s-1}$ possibilities. Denote by $\bm{q}^{k}=[q_{0}^{k},\ldots,q_{s-2}^{k}]$ the $k$:th possibility, meaning that $q_{j}^{k}$ is either $r_{j}$ or $1/\overline{r_{j}}$ for $20\leq j\leq s-$ 2, $10\leq k\leq 2^{s-1}-$ 1. For each $\bm{q}^{k}$, we get an equation system in $\bm{g}$
@@
The equation system in (109) with $s-1$ equations represents the zeros of a polynomial of degree $s-1$ and thus it has a unique solution $\bm{g}^{k}$ up to a complex scalar. Each $\bm{g}^{k}$, with proper normalization, is a solution to (72), i.e., $\bm{y}^{+}=|\bm{V}^{T}(\bm{z})\bm{V}(\bm{\theta}^{+})\bm{g}^{k}|^{2}$. Therefore, after $14s-$ 1 measurements as in R4 (with shifted harmonic samples) there are $2^{s-1}$ solutions to R4. To resolve the ambiguity of which $\bm{g}^{k}$ equals $\bm{g}^{+}$, we need an additional measurement - this is the purpose of the arbitrary vector $\bm{a}$ in R5. W.p.1, only one of $\bm{g}^{k}$, $10\leq k\leq 2^{s-1}-$ 1, will satisfy $y_{m}^{+}=|\bm{a}^{T}\bm{V}(\bm{\theta}^{+})\bm{g}^{k}|^{2}$, resolving the ambiguity and thus giving us $\bm{g}^{+}$.
In the derivations above, we assumed that $g_{k}\neq 0$, $10\leq k\leq s-$ 1; i.e., we assumed exact knowledge of the number of non-zero elements. If $g_{k}=0$ for some $k$, $\tilde{\bm{G}}(\bm{z})$ in (III-A) can have a null space of more than one dimension. Assume that $S$, $S\leq s$, of the $g_{0},\ldots,g_{s-1}$ are non-zero. One can now iterate through different realizations of $\tilde{\bm{G}}(\bm{z})$, reducing the value of $s$ in each step by 1 until a one-dimensional null space is encountered. Since for each such realization of $\tilde{\bm{G}}(\bm{z})$, $\bm{w}^{+}$ in (83) belongs to its null space, the solution will not be missed and the iteration terminates after at most $s-S+1$ steps (when $s$ is reduced to $S$, giving the unique solution $c\bm{w}^{+}$ as implied by the derivations above).
@@ 3. Output: $\bm{\theta}^{+},e^{i\alpha}\bm{g}^{+}$ for some unknown $\alpha$
    Construct the Laurent polynomial $\hat{L}(z)=\sum_{k=0}^{2S}w_{k}z^{k-S}$. The conjugate of the unique roots of $\hat{L}(z)$ equal $\bm{\theta}^{+}$. $|\bm{g}^{+}|$ can now be computed (up to a global real scalar) from Proposition 5.
@@
  From the theoretical results above, it follows that R5-Alg-Harmonic recovers $\bm{\theta}^{+}$ and $\bm{g}^{+}$ w.p.1. Steps 1-6 produce all solutions to R4 while step 7 finds the correct $\bm{g}^{+}$ and thereby gives the result in R5. The complexity of R5-Alg-Harmonic is $O(\mathrm{poly}(s)2^{s})$. Up until step 6, the total complexity is $O(\mathrm{poly}(s))$. Solving (109) has complexity $O(\mathrm{poly}(s))$ but it is repeated $2^{s-1}$ times in step 6. It is possible to utilize Theorem 6 to directly generate the dual solution for each $\bm{g}^{k}$ and thereby reduce the number of times (109) is solved by half, but this still results in complexity $O(\mathrm{poly}(s)2^{s})$. Fortunately, it turns out that by taking $24s-$ 2 additional measurements, the number of possible solutions can be reduced from $2^{s-1}$ to only two.
@@
Given the assumptions in Theorem 7, further assume that $z_{j}^{n}=e^{i\omega}$, $\hat{L}^{+}(z_{j})\neq 0$ for $34s-1\leq j\leq 6s-$ 3 and some $\omega$ satisfying $e^{i\gamma}\neq e^{i\omega}$ (with $\gamma$ from Theorem 7); $z_{j}^{n}=e^{i\phi}$, $\hat{L}^{+}(z_{j})\neq 0$ for $46s-2\leq j\leq 8s-$ 4 and some $\phi$ satisfying $e^{i\phi}\neq e^{i\gamma}$ and $(e^{i\phi}-e^{i\gamma})(e^{-i\omega}-e^{-i\gamma})\neq(e^{-i\phi}-e^{-i\gamma})$. Solving (60) (or equivalently (III-A)) results in $\hat{L}(z)=c\hat{L}^{+}(z),L(z)=cL^{+}(z),\tilde{L}(z)=c\tilde{L}^{+}(z)$ for some real scalar $c$.
@@
The assumptions on $\phi,\omega,\gamma$ imply that the scalar in front of $\tilde{L}(z)-c\tilde{L}^{+}(z)$ in (117) is non-zero. Thus, the identity theorem for polynomials implies that (117) given $\tilde{L}(z)=c\tilde{L}^{+}(z)$. Inserting this relation into (110) we also obtain $L(z)=cL^{+}(z)$, finishing the proof. ∎
Clearly, Theorem 8 holds even if $m\geq 8s-3$ as long as a subset of $38s-$ 3 measurements satisfies the assumptions of Theorem 8. Arbitrary $\bm{\theta}^{+}\in\mathbb{T}^{s}$ and $\bm{g}^{+}$ satisfy the assumptions of Theorem 8 w.p.1, so that $\bm{G}(\bm{z})$ in (III-A) has a one-dimensional null space w.p.1. Moreover, for any $\bm{\theta}^{+}\in\mathbb{T}^{s}$ and $\bm{g}^{+}$ with no zero elements, it is evidently easy to choose a $\bm{z}$ satisfying the assumptions of Theorem 8. Similarly to Theorem 5, we can relax the requirements on $\bm{z}$ from Theorem 8.
@@
Assume that $n\geq 4s-1$. Given any $\bm{\theta}^{+}\in\mathbb{T}^{s}$ and any $\bm{g}^{+}\in\mathbb{C}^{s}$ with $s$ non-zero elements, solving (60) (or equivalently (III-A)) for an arbitrary $\bm{z}\in\mathbb{T}^{8s-3}$ results in $\hat{L}(z)=c\hat{L}^{+}(z),L(z)=cL^{+}(z),\tilde{L}(z)=c\tilde{L}^{+}(z)$ (for some real scalar $c$) w.p.1.

### 06 Feb 2024 v2

@@ II-B Derivation of R2
    Compute $\bm{g}=\left(\bm{V}(\bm{z})^{T}\bm{V}(\bm{\theta}_{\bm{p}})\right)^{\dagger}\bm{y}$ or use the procedures outlined in the proofs of Corollary 1 and Corollary 2.
@@ III Phase-less Measurements
  R3: There exists a construction of an $m\times n$ measurement matrix $\bm{A}$, with $m\geq 4s$ and $n\geq 4s-1$, such that for almost all $\bm{x}\in\mathbb{C}_{\leq s}^{n}$ the solution to
$\bm{y}=\left|\bm{A}\bm{x}\right|^{2}$ (52)
is unique and therefore equals $\bm{x}$ (up to a global phase). Furthermore, there exists a corresponding recovery algorithm such that
@@
  This result provides a solution to problem 3.5 in [18] for almost all $\bm{x}\in\mathbb{C}_{\leq s}^{n}$. Moreover, the result is significantly better than asked for by 3.5 since R3 achieves $O(\mathrm{poly}(s))$ recovery complexity with $O(s)$ measurements. To obtain R3, we begin by considering the phase-less equivalent of R1 with a Fourier matrix
@@
  • If $4s-1\leq m<\min\{n+1,8s-3\}$ and samples in $\bm{z}$ are shifted harmonics, $\bm{\theta}$ and $|\bm{g}|$ (up to a global real positive scalar) can be exactly recovered with $O(\mathrm{poly}(s))$ complexity and there are $2^{s-1}$ different $\bm{g}$ satisfying (53).
  • If $m\geq 8s-3$ and $\bm{z}$ is arbitrary, $\bm{\theta}$ and $|\bm{g}|$ (up to a global real positive scalar) can be exactly recovered with $O(\mathrm{poly}(s))$ complexity and there are only two different $\bm{g}$ satisfying (53).
To derive R4 we will start (as with R1) by first recovering $\bm{\theta}$ from (53). However, in contrast to R1, we can no longer recover $\bm{g}$ with a simple matrix inversion due to the non-linear magnitude operator. After support recovery $\bm{\theta}$, we are (at first glance) essentially back to a classical phase retrieval problem. Below we will present an approach, similar in nature to the approach in Section II, from which certain Laurent polynomials are solved for. It turns out that there is enough information about $\bm{g}$ embedded in these Laurent polynomials for recovering $\bm{g}$.
R4 shows that there is an ambiguity when measuring as in (53) which cannot be resolved by simply taking more measurements (increasing $m$). In order to resolve the ambiguity of which $\bm{g}$ is the correct one, one typically utilizes side information about the signal (e.g. knowledge of one element in $\bm{g}$ or $\bm{x}$) as explained in [6] or takes one additional measurement with some arbitrary measurement vector $\bm{a}$. Therefore, we have the following reformulation of R4
@@
  • If $m\geq 8s-3$ and for almost all $\bm{z}$, $\bm{\theta}$ and $\bm{g}$ can be recovered exactly with $O(\mathrm{poly}(s))$ complexity.
@@ III-A Derivation of R4 and R5
where the operator ${.}\mathrm{fliplr}\{.\}$ flips a matrix vertically (i.e., flips its columns), $\bm{G}(\bm{z})$ has dimensions $m\times(8s-2)$ and $\bm{w}=[\hat{\bm{l}};\tilde{\bm{l}};\bm{l};\overline{\tilde{\bm{l}}}]$. Assume that $S$ values from $g_{0}^{+},\ldots,g_{s-1}^{+}$ are non-zero. Since $\hat{L}(z)=\hat{L}^{+}(z),L(z)=L^{+}(z),\tilde{L}(z)=\tilde{L}^{+}(z)$ solves (60),
@@
solves (III-A) and vice versa. This implies that $\bm{G}(\bm{z})$ cannot be full rank and therefore has a null space of at least one dimension. If $\bm{G}(\bm{z})$ has a one-dimensional null space, meaning that (III-A) has the solution $\bm{w}=c\bm{w}^{+}$ for any real positive scalar $c$, then (60) has the unique solution $\hat{L}(z)=c\hat{L}^{+}(z),L(z)=cL^{+}(z),\tilde{L}(z)=c\tilde{L}^{+}(z)$ and vice versa.
@@
for any real positive scalar $c$, which is equivalent to the solution $\hat{L}(z)=c\hat{L}^{+}(z)$, $L(z)+e^{i\gamma/n}\tilde{L}(z)+e^{-i\gamma/n}\overline{\tilde{L}(z)}=c\left(L^{+}(z)+e^{i\gamma/n}\tilde{L}^{+}(z)+e^{-i\gamma/n}\overline{\tilde{L}(z)^{+}}\right)$ of (72) and vice versa.
@@
Assume that $n\geq 4s-1$, $z_{j}^{n}=e^{i\gamma}$ for $20\leq j\leq 4s-$ 2, $g_{k}^{+}\neq 0$, $\theta_{k}^{+}\in\mathbb{T}$ and $(\theta_{k}^{+})^{n}\neq e^{-i\gamma}$ for $10\leq k\leq s-$ 1. Solving (72) (or equivalently (III-A)) results in $\hat{L}(z)=c\hat{L}^{+}(z),L(z)+z^{n}\tilde{L}(z)+z^{-n}\overline{\tilde{L}}(z)=c\left(L^{+}(z)+z^{n}\tilde{L}^{+}(z)+z^{-n}\overline{\tilde{L}^{+}}(z)\right)$ for some real positive scalar $c$.
@@
Since $(\theta_{k}^{+})^{n}\neq e^{-i\gamma}$ and $g_{k}\neq 0$, it follows that $t_{k}^{+}\left(\frac{1}{\theta_{k}^{+}}\right)\neq 0$ and thus $g_{k}^{+}t_{k}^{+}\left(\frac{1}{\theta_{k}^{+}}\right)\neq 0$, which implies that the roots of $\hat{L}^{+}(z)$ are not roots of $L^{+}(z)+e^{i\gamma}\tilde{L}^{+}(z)+e^{-i\gamma}\overline{\tilde{L}^{+}(z)}$, a contradiction. Hence, the roots of $\hat{L}^{+}(z)$ are roots of $\hat{L}(z)$ which implies that $\hat{L}(z)=c\hat{L}^{+}(z)$ for some real positive scalar $c$.
@@
From the result in Theorem 7, it follows that when $g_{k}^{+}\neq 0$, $10\leq k\leq s-$ 1, we can find $\hat{L}^{+}(z)$ (up to a real positive scalar $c$) after $14s-$ 1 measurements with shifted harmonic samples from which we can recover the support $\bm{\theta}^{+}$ (below we will deal with the case when $S$, $S\leq s$, of $g_{0}^{+},\ldots,g_{s-1}^{+}$ are non-zero). Furthermore, from these measurements, we also obtain $L^{+}(z)+e^{i\gamma}\tilde{L}^{+}(z)+e^{-i\gamma}\overline{\tilde{L}^{+}(z)}=\left|e^{i\gamma}\hat{u}^{+}(z)+\tilde{u}^{+}(z)\right|^{2}$ (up to a real positive scalar $c$). Given samples that are shifted harmonics, almost all $\bm{\theta}^{+}\in\mathbb{T}^{s}$ and $\bm{g}^{+}\in\mathbb{C}^{s}$ (with $s$ non-zero elements) satisfy the assumptions of Theorem 7. Similarly, for any $\bm{\theta}^{+}\in\mathbb{T}^{s}$ and $\bm{g}^{+}\in\mathbb{C}^{s}$ (with $s$ non-zero elements), almost all $\gamma$-shifted harmonics satisfy the assumptions of Theorem 7.
From Theorem 7 one can find $|\bm{g}^{+}|$ up to a real positive scalar.
Corollary 3.
Given the realizations $\bm{\theta}^{+}$, $\bm{g}^{+}$ of $\bm{\theta}$, $\bm{g}$, respectively. For almost all vectors $\bm{z}$ that consist of at least $14s-$ 1 shifted harmonics, $\bm{\theta}^{+}$ and $|\bm{g}^{+}|$ can be found, up to a real positive scalar, with $O(\mathrm{poly}(s))$ complexity.
@@
for some unknown real positive scalar $c$. Since $(\theta_{k}^{+})^{-1}=\overline{\theta_{k}^{+}}$ and
@@
Clearly, the operations leading to $c|\bm{g}^{+}|^{2}$ are of $O(\mathrm{poly}(s))$ complexity, showing that $\sqrt{c}|\bm{g}^{+}|$ can be found with $O(\mathrm{poly}(s))$ complexity. ∎
@@
When it comes to recovering $\bm{g}^{+}$, the situation is trickier. Given $\left|e^{i\gamma}\hat{u}^{+}(z)+\tilde{u}^{+}(z)\right|^{2}$, up to a real positive scalar, Proposition 2 implies that the roots of $\left|e^{i\gamma}\hat{u}^{+}(z)+\tilde{u}^{+}(z)\right|^{2}$ appear as conjugate pairs: if $r$ is a root of $\left|e^{i\gamma}\hat{u}^{+}(z)+\tilde{u}^{+}(z)\right|^{2}$ then so is $1/\overline{r}$ as well. Let $(r_{0},1/\overline{r_{0}}),\ldots,(r_{s-2},1/\overline{r_{s-2}})$ be the $s-1$ conjugate root pairs. As seen from (II-A), $e^{i\gamma}\hat{u}^{+}(z)+\tilde{u}^{+}(z)$ is of the same form as $\tilde{u}^{+}(z)$ in (II-A) but with coefficients $(e^{i\gamma}(\theta_{k}^{+})^{n}-1)g_{k}^{+}$, $10\leq k\leq s-$ 1. Hence, for almost all $\bm{g}^{+}$, Lemma 2 implies that $(r_{0},1/\overline{r_{0}}),\ldots,(r_{s-2},1/\overline{r_{s-2}})$ are distinct. Only one member from each pair can be a root of $e^{i\gamma}\hat{u}^{+}(z)+\tilde{u}^{+}(z)$. Since we do not know beforehand which member of each pair is the actual root of $e^{i\gamma}\hat{u}^{+}(z)+\tilde{u}^{+}(z)$, we need to go through all the $2^{s-1}$ possibilities. Denote by $\bm{q}^{k}=[q_{0}^{k},\ldots,q_{s-2}^{k}]$ the $k$:th possibility, meaning that $q_{j}^{k}$ is either $r_{j}$ or $1/\overline{r_{j}}$ for $20\leq j\leq s-$ 2, $10\leq k\leq 2^{s-1}-$ 1. For each $\bm{q}^{k}$, we get an equation system in $\bm{g}$
@@
The equation system in (109) with $s-1$ equations represents the zeros of a polynomial of degree $s-1$ and thus it has a unique solution $\bm{g}^{k}$ up to a complex scalar. Each $\bm{g}^{k}$, with proper normalization, is a solution to (72), i.e., $\bm{y}^{+}=|\bm{V}^{T}(\bm{z})\bm{V}(\bm{\theta}^{+})\bm{g}^{k}|^{2}$. Therefore, after $14s-$ 1 measurements as in R4 (with shifted harmonic samples) there are $2^{s-1}$ solutions to R4. To resolve the ambiguity of which $\bm{g}^{k}$ equals $\bm{g}^{+}$, we need an additional measurement - this is the purpose of the arbitrary vector $\bm{a}$ in R5. With probability 1, only one of $\bm{g}^{k}$, $10\leq k\leq 2^{s-1}-$ 1, will satisfy $y_{m}^{+}=|\bm{a}^{T}\bm{V}(\bm{\theta}^{+})\bm{g}^{k}|^{2}$, resolving the ambiguity and thus giving us $\bm{g}^{+}$.
In the derivations above, we assumed that $g_{k}^{+}\neq 0$, $10\leq k\leq s-$ 1; i.e., we assumed exact knowledge of the number of non-zero elements. If $g_{k}^{+}=0$ for some $k$, $\tilde{\bm{G}}(\bm{z})$ in (III-A) can have a null space of more than one dimension. Assume that $S$, $S\leq s$, of the $g_{0}^{+},\ldots,g_{s-1}^{+}$ are non-zero. One can now iterate through different realizations of $\tilde{\bm{G}}(\bm{z})$, reducing the value of $s$ in each step by 1 until a one-dimensional null space is encountered. Since for each such realization of $\tilde{\bm{G}}(\bm{z})$, $\bm{w}^{+}$ in (83) belongs to its null space, the solution will not be missed and the iteration terminates after at most $s-S+1$ steps (when $s$ is reduced to $S$, giving the unique solution $c\bm{w}^{+}$ as implied by the derivations above).
@@ 3. Output: $\bm{\theta}^{+},e^{i\alpha}\bm{g}^{+}$ for some unknown $\alpha$
    Construct the Laurent polynomial $\hat{L}(z)=\sum_{k=0}^{2S}w_{k}z^{k-S}$. The conjugate of the unique roots of $\hat{L}(z)$ equal $\bm{\theta}^{+}$. $|\bm{g}^{+}|$ can now be computed (up to a global real positive scalar) from Corollary 3.
@@
  From the theoretical results above, it follows that R5-Alg-Harmonic recovers almost all $\bm{\theta}^{+}$ and $\bm{g}^{+}$. Steps 1-6 produce all solutions to R4 while step 7 finds the correct $\bm{g}^{+}$ and thereby gives the result in R5. The complexity of R5-Alg-Harmonic is $O(\mathrm{poly}(s)2^{s})$. Up until step 6, the total complexity is $O(\mathrm{poly}(s))$. Solving (109) has complexity $O(\mathrm{poly}(s))$ but it is repeated $2^{s-1}$ times in step 6. It is possible to utilize Theorem 6 to directly generate the dual solution for each $\bm{g}^{k}$ and thereby reduce the number of times (109) is solved by half, but this still results in complexity $O(\mathrm{poly}(s)2^{s})$. Fortunately, it turns out that by taking $24s-$ 2 additional measurements, the number of possible solutions can be reduced from $2^{s-1}$ to only two.
@@
Given the assumptions in Theorem 7, further assume that $z_{j}^{n}=e^{i\omega}$, $\hat{L}^{+}(z_{j})\neq 0$ for $34s-1\leq j\leq 6s-$ 3 and some $\omega$ satisfying $e^{i\gamma}\neq e^{i\omega}$ (with $\gamma$ from Theorem 7); $z_{j}^{n}=e^{i\phi}$, $\hat{L}^{+}(z_{j})\neq 0$ for $46s-2\leq j\leq 8s-$ 4 and some $\phi$ satisfying $e^{i\phi}\neq e^{i\gamma}$ and $(e^{i\phi}-e^{i\gamma})(e^{-i\omega}-e^{-i\gamma})\neq(e^{-i\phi}-e^{-i\gamma})$. Solving (60) (or equivalently (III-A)) results in $\hat{L}(z)=c\hat{L}^{+}(z),L(z)=cL^{+}(z),\tilde{L}(z)=c\tilde{L}^{+}(z)$ for some real positive scalar $c$.
@@
The assumptions on $\phi,\omega,\gamma$ imply that the scalar in front of $\tilde{L}(z)-c\tilde{L}^{+}(z)$ in (117) is non-zero. Thus, the identity theorem for polynomials implies that $\tilde{L}(z)=c\tilde{L}^{+}(z)$ must hold in (117). Inserting this relation into (110) we also obtain $L(z)=cL^{+}(z)$, finishing the proof. ∎
Clearly, Theorem 8 holds even if $m\geq 8s-3$ as long as a subset of $38s-$ 3 measurements satisfies the assumptions of Theorem 8. For any $\bm{z}$ consisting of shifted harmonics, with the different shifts specified in Theorem 8, almost $\bm{\theta}^{+}\in\mathbb{T}^{s}$ and $\bm{g}^{+}$ satisfy the assumptions of Theorem 8. Similarly, for any given $\bm{\theta}^{+}\in\mathbb{T}^{s}$ and $\bm{g}^{+}$ with no zero elements, almost all $\bm{z}$ consisting of shifted harmonics as specified in Theorem 8 satisfy the assumptions of Theorem 8. Similar to Theorem 5, we can relax the assumptions on $\bm{z}$.
@@
Assume that $n\geq 4s-1$ and $m=8s-3$. Given any $\bm{\theta}^{+}\in\mathbb{T}^{s}$ and any $\bm{g}^{+}\in\mathbb{C}^{s}$ (with $s$ non-zero elements), solving (60) (or equivalently (III-A)) results in $\hat{L}(z)=c\hat{L}^{+}(z),L(z)=cL^{+}(z),\tilde{L}(z)=c\tilde{L}^{+}(z)$ (for some real positive scalar $c$) for almost all $\bm{z}\in\mathbb{T}^{m}$. Similarly, given almost any $\bm{z}\in\mathbb{T}^{m}$, solving (60) (or equivalently (III-A)) gives that $\hat{L}(z)=c\hat{L}^{+}(z),L(z)=cL^{+}(z),\tilde{L}(z)=c\tilde{L}^{+}(z)$ for almost all $\bm{\theta}^{+}\in\mathbb{T}^{s}$ and $\bm{g}^{+}\in\mathbb{C}^{s}$ (with $s$ non-zero elements).